How much would you invest today in order to receive $50,000 in each of the following scenarios?

(Click here to see present value and future value tables)

Round your answers to 2 decimal places.

A. 11 years at 10% $
B. 7 years at 12% $
C. 14 years at 15% $
D. 20 years at 20% $
Answer #1

A. Invest today = Amount x Present value Factor

= 50000*0.350 = $17,500

B. Invest today = 50000*0.452 = $22,600

C. Invest today = 50000*0.141 = $7,050

D. Invest today = 50000*0.026 = $1,300